OpenStudy (anonymous):

Find the GCF of 200 and 75. A. 15 B. 25 C. 50 D. 75 @some.random.cool.kid @geerky42

OpenStudy (anonymous):

25

OpenStudy (anonymous):

b

OpenStudy (anonymous):

ok

geerky42 (geerky42):

200 = 25 * 8 75 = 25 * 3

geerky42 (geerky42):

So yeah it's B.

OpenStudy (anonymous):

ok thx